OpenStudy (anonymous):

solve the equation below for x: log2x+log(2x-1)=log3x

OpenStudy (anonymous):

mmmhhh i dont really know how to start this process ???

OpenStudy (ikram002p):

base is 10 for all ?

OpenStudy (ikram002p):

if yes then hint log(a.b)=log a + log b

OpenStudy (anonymous):

it does not specified anything like that @ikram002p

OpenStudy (anonymous):

Don't worry about base here

OpenStudy (ikram002p):

ok so first term is log (2x) not \(\log_2 x\) right ?

OpenStudy (ikram002p):

then use the hint i gave ^^

OpenStudy (anonymous):

\[\log2x+\log(2x-1)=\log3x\] thats the equation

OpenStudy (ikram002p):

log2x+log(2x-1)=log3x log(a.b)=log a + log b log[2x(2x-1)]=log3x

OpenStudy (ikram002p):

ok i git u nw @meln_n ^^ can u use the Hint ?

OpenStudy (anonymous):

mmmhhh im trying hehhe lets seee

OpenStudy (anonymous):

log2x+log(2x-1)=log3x log 2x*(2x-1) = log 9x 2x*(2x-1) = 3x Now solve this for x.

OpenStudy (ikram002p):

yeah check @ShailKumar work ^

OpenStudy (anonymous):